THE PFI-ZEKE SPECTROSCOPY STUDY OF HfS^{+} AND THE IONIZATION ENERGY OF HfS

\underline{I. O. ANTONOV}, B. J. BARKER, M. C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Spectroscopic data for the ground and low-lying states HfS^{+} have been obtained using the technique of pulse field ionization - zero electron kinetic energy (PFI-ZEKE) spectroscopy. PFI-ZEKE spectra were recorded for the levels X^{2}Σ^{+} (v=0-18), 2Δ_{5/2} (v=0-8) and ^{2}Δ_{3/2} (v=0-3). Assignments of the electronically excited states of HfS^{+} are based on CCSD(T) and DFT calculations with SDB-aug-cc-pVTZ basis set. Rotationally resolved spectra were recorded for the X^{2}Σ^{+} (v=0) state using single rotational line excitation of the intermediate state. The ionization energy for HfS, term energies and molecular constants for the ground and low-lying states of HfS^{+} will be reported.